### v2

# Genixer: Empowering Multimodal Large Language Model as a Powerful Data Generator

Abstract
Instruction tuning data is essential for training the Multimodal Large Language Models (MLLMs). However, the creation of high-quality instruction tuning data presents significant challenges. Prior methods that depended on GPT-4 for data generation were not only costly but also lacked satisfactory performance in complex tasks (i.e., grounding-based reasoning tasks). To address these issues, we developed an innovative data generation pipeline Genixer to generate various high-quality instruction tuning data, including nine representative tasks, e.g., Common VQA, REC, REG, and PointQ. Specifically, Genixer provides a unified solution with four key steps for alleviating the difficulty of data generation: (i) instruction data collection, (ii) instruction template design, (iii) empowering MLLM, and (iv) data generation and filtering. Subsequently, the superior qualitative results of our Genixer demonstrate that current MLLMs have a strong potential to evolve into powerful data generators. Additionally, to validate the efficacy of generated data quantitatively, we add the instruction tuning data produced by Genixer into the training of two representative MLLMs and observe the consistent improvements on various VQA tasks and multimodal benchmarks. The data, code, and models will be released.
Keywords:
Multimodal Large Language ModelInstruction Tuing1 Introduction
Large Language Models (LLMs) [6, 37, 2, 14] have achieved remarkable success in understanding human instructions and demonstrated their high proficiency in tackling complex natural language tasks. This progress has recently extended to Multimodal Large Language Models (MLLMs) [17, 40, 13, 29, 53, 14, 23, 9, 51, 45, 55, 3], where visual information is encoded by a dedicated visual encoder and seamlessly integrated into LLMs. This integration empowers MLLMs with the ability to analyze and reason about multimodal tasks effectively.
The training of a satisfactory MLLM heavily relies on the availability of sufficient high-quality visual instruction tuning data. The data are, in many ways, the lifeblood of contemporary general-purpose intelligence systems, much like oil powers industrial processes. To attain superior performance, one pioneer model like InstructBLIP [13] requires training on multiple held-in datasets, such as VQAv2 [15] and OKVQA [35], from traditional Vision-Language (VL) tasks. Unfortunately, these held-in datasets suffer from a limitation in image diversity, as most of them originate from the COCO dataset [27], potentially restricting the model generalization ability on unseen visual data. This naturally raises the fundamental question: how can we obtain a sufficient quantity of high-quality instruction tuning data for training MLLMs?
As of now, there is no effective solution to this challenge yet. The traditional approach of manual labeling proves prohibitively expensive and impractical, particularly since MLLM training often necessitates million-scale instruction tuning data to ensure robust generalization. Another potential solution is to leverage powerful models to generate high-quality data. Unfortunately, almost all open-sourced pretrained MLLMs fall short in this regard. As exemplified in Fig. 1, models like LLaVA-1.5 [28], mPLUG-owl [51], and Shikra [9] tend to treat conversation generation as a dense captioning task and struggle to produce natural and coherent dialog data and IntructBLIP [13] directly fails. Contrarily, our produced dialogue exhibits both versatility and intelligence, attaining a level comparable to that of GPT-4V [38]. Except for leveraging current MLLMs, another viable way is to utilize GPT-4V with carefully designed prompts via the released API, but such a choice leads to high financial costs. For instance, producing millions of samples, each with thousands of tokens, with GPT-4V [38] would cost millions. Moreover, as illustrated in Fig. 2, GPT-4V cannot generate satisfied instruction tuning data of some complex data types (e.g., referring expression comprehension), further limiting its usage for exclusive model learning.
In this work, we address the challenge of generating highly qualified multimodal data. To this end, we introduce Genixer, an automatic data generation pipeline that provides a unified solution with four key steps: (i) instruction data collection, (ii) instruction template design, (iii) empowering MLLM, and (iv) data generation and filtering. We begin by categorizing current Vision-Language (VL) tasks into general and grounding tasks. These categories include nine data types, e.g., Common VQA, Multi-choice VQA, Referring Expression Comprehension (REC), and Referential Dialogue (RD). See details in Sec. 3.1. Then, we compile datasets corresponding to these nine tasks and consistently reformat the input data into the instruction-following format. To acquire the flexibility of automatic large-scale and user-guided data generation, we propose a two-level instruction mechanism for creating data in two modes: type-agnostic generation and type-specific generation (see Sec. 3.2). In the type-agnostic mode, given one image, users can pose an instruction without specifying a particular data type to allow the model to perform data generation freely. This is because, in some cases, preemptively determining the optimal data type for generation becomes particularly daunting, especially when it involves large-scale data production for model training purposes. For the type-specific mode, users can ask the model to generate a specific data type from the above nine data types. Subsequently, we leverage two representative models, LLaVA1.5 [28] and Shikra [9] evolve them into the intelligent data generators with the abovementioned two-level instruction template yielding two data generators, Genixer and Genixer, for general and grounding data generation, respectively (See Sec. 3.3). To maintain high data quality, we employ the third-party MLLM Fuyu-8B [5] as an evaluator within our specially designed evaluation framework for general task evaluation, detailed in Sec. 3.4. Additionally, we utilize CLIP [19] to calculate the region-expression similarity for REC-like data, enabling us to filter out samples exhibiting low similarity.
Experimental results show that using 915K VQA-like tuning data generated by Genixer to train LLaVA1.5 can make a significant performance enhancement compared to the vanilla LLaVA1.5 across a diverse set of tasks, such as 3.8% improvement on Vizwiz [16] and 2.7% on SicenceQA [31]. Furthermore, our Genixer extends its utility to generate instruction-tuning data for visual grounding tasks akin to those encountered in the SoTA grounding model, Shikra. These results underscore the remarkable efficacy of our proposed data generation pipeline, Genixer. In summary, our contributions are:
We introduce an innovative multimodal data generation pipeline, Genixer, that enables existing MLLM models to function as potent data generators to effectively and affordably generate high-quality instruction data.
We contribute two data generators Genixer and Genixer for a wide range of multimodal instruction tuning data generation.
We contribute two high-quality multimodal data: 915K VQA-like data, which can greatly improve LLaVA1.5 on various tasks like Vizwiz and ScienceQA, and 350K REC-like data that enhance Shikra on the grounding tasks like REC datasets.
2 Related Work
Multimodal Large Language Models. Large Language Models (LLMs) showcased the remarkable complicated reasoning abilities. Some well-known open-souced LLMs include FlanT5 [12], OPT [54], LLaMA [46], Vicuna [11] and LLaMA-2 [47] show exceptional reasoning ability of solving math, codes problems. By leveraging these LLMs, a surge of multimodal modes [25, 29, 13, 3, 9, 49, 55, 8, 7, 32, 50, 17, 40, 36, 51, 14] are proposed to integrate the visual information for diverse multimodal reasoning tasks such as image captioning [10, 1, 52] and visual question answering [15, 35, 18, 16]. LLaVA [29] is a pioneering approach that adopts a single linear layer to project the CLIP [42] encoder extracted visual features to the input of LLM. Different from LLaVA, InstructBLIP [13] employs an instruction-aware feature extractor and obtains advanced performance on various tasks building upon the pretrained BLIP2 [25]. Besides focusing on traditional multimodal tasks, Shikra [9] and PVIT [7] pay attention to the grounding multimodal tasks by carefully designing the visual instruction data or employing the region-based vision encoder. Except for these MLLMs, CogVLM [49], Qwen-VL [3], and Kosmos-2 [40] explore adopting a billion-scale pertaining data corpus to enhance the model generalization ability and robustness.
Multimodal Instruction Data. High-quality multimodal instruction data is crucial for training a robust MLLM. Two main setups are as follows: 1) Transforming the current vision-language datasets into the instruction-tuning format, e.g., InstructBLIP. Such choice is limited by the diversity of image sources. 2) Some approaches LLaVA, VisionLLM[50], and Shikra, resort to prompting the GPT-4 [37] language model to generate corresponding instruction data. This way requires the image datasets to have enough captions or region-based annotations (e.g., bounding boxes), which heavily restricts the data scale. Additionally, prompting commercial LLMs incurs high costs, and even GPT-4V [38] may not address the data generation effectively on some specific tasks, as illustrated in Fig. 2. To this end, we introduce Genixer, an innovative pipeline that explores the capabilities of MLLMs to generate high-quality multimodal instruction data.
3 Genixer: An Automatic Instruction Tuning Data Generation Pipeline
Category | Task | Dataset | Size |
General | Common VQA | VQAv2, GQA, Counting110K, POPE | 583K |
Adv VQA | POPE | 20K | |
MC VQA | A-OKVQA | 17K | |
MD | VQAv2, LLaVA-Conv-58K | 108K | |
Grounding | REC | VG, RefCOCO | 1M |
REG | VG, RefCOCO | 1M | |
PointQA | PointQA Local, Visual7W | 218K | |
QCA | Shikra (GPT-4 Generated) | 4K | |
RD | Shikra (GPT-4 Generated) | 1.8K |
Though current MLLMs show exceptional capability of handling various multimodal tasks [28, 3, 24], rare works are concentrating on exploring the data generation capability. In this work, instead of focusing on how to make an MLLM a good problem-solver, we try to provoke the potential ability of current MLLM to generate instruction tuning data automatically and intelligently. To this end, We propose Genixer, as illustrated in Fig. 3, which is a novel pipeline that contains four key steps, including 1) instruction data collection, 2) instruction template design, 3) empowering MLLMs, and 4) data generation and filtering. In the following, we will elaborate on these four key steps.
3.1 Instruction Data Collection
In accordance with the prevalence and practical relevance of real-world multimodal tasks, we carefully select nine representative multimodal tasks as listed in Tab. 1 for training Genixer, including Common Visual Question Answering (Common VQA), Adversarial-based VQA (Adv VQA), Multi-choice VQA (MC VQA), Multi-turn Dialogue (MD), Referring Expression Comprehension (REC), Referring Expression Generation (REG), PointQA, QCA and Referential Dialogue (RD). We divide these tasks into two categories, general tasks and grounding tasks, and display some examples in Fig. 4 for visual inspection.
3.2 Instruction Template Design
In an automatic data generation context, where image content is agnostic, preemptively determining the optimal data type for generation becomes particularly daunting, especially when it involves large-scale data production for model training purposes. For instance, in a landscape picture without any humans or objects, it is hard to ask a high-quality multi-choice question, and it is easy to generate an adversarial question such as “Is there a boat in the image?” Hence, it is imperative to delineate two key modes for a robust data-generative MLLM: 1) type-agnostic data generation and 2) type-specific data generation.
Two-level Instructions. To achieve this goal, we propose a two-level instruction template for achieving controllable question-answer generation at the inference stage. Specifically, the complete instruction design is as follows:
The tags <image>, <question>, and <answer> serve as placeholders for inserting the tokens of the image, question, and answer, respectively. Question: <question> Answer: <answer> is the model response that needs to be predicted in a left-to-right text generation manner.
Regarding General Instruction, we meticulously compile a list of instructions, and for each piece of training data, we randomly choose one example to use during the training process. For instance, “Please provide a clear and direct question and answer after examining the image”. This instruction plays a role in allowing the model to generate diverse types of instruction tuning data without any specific condition referred to as mode 1.
Then, for Specific Instruction, it denotes designating the specific question type, such as “This is a Common VQA task”, allowing us to control the output type of the generated data referred to as mode 2.
Controlling Constant. During the training phase, we set a constant for controlling the ratio of training samples that are exclusive with General Instruction. Consequently, in the inference phase, the model is able to switch the mode by adding specific instructions or not. For example, as illustrated in Fig. 5, the model is capable of generating various types of data in the absence of specific instructions. Simultaneously, it can produce specific outputs when guided by a detailed prompt, like “This is an MC VQA task”.
3.3 Empowering Current MLLMs
A robust model with multimodal reasoning ability is the cornerstone for building an instruction tuning data generator. Therefore, we resort to delineating the current remarkable MLLM as the base model and then fine-tuning it with our developed two-level instruction template. Consequently, we present two data generators, Genixer and Genixer, to respectively generate instruction tuning data for the two data categories, general tasks, and grounding tasks as summarized in Tab. 1.
Overall Framework of Genixer and Genixer. For brevity, we denote the MLLM model as , the input general and specific instructions as and , respectively. Then, given an image , our target is to enforce to generate question and corresponding answer :
To this end, we follow previous MLLMs[28, 13], and design the training objective in an autoregressive manner:
where is the whole sentence compose and , and is current prediction token. is the length of the model response sequence. denotes the total trainable parameters in (e.g., the parameters of projector and LLM with LLaVA1.5 backbone). In the following, we select LLaVA1.5 and Shikra to generate two kinds of data because of their SoTA performance in the corresponding areas. It is noteworthy that Genixer is flexible and applicable to work with any strong MLLM.
Training of Genixer. Genixer trains the pretrained LLaVA1.5 [28], a SoTA model for four kinds of general tasks, including Common VQA, Adv VQA, MC VQA, and MD. As summarized in Tab. 1, we only sample a subset of data for these four data types to improve training efficiency. The set to 0.2, 0.2, 0.5, 0.2, respectively. The different ratios are because of the different data sizes of these four data types. To keep the training data balanced, we handcrafted the values. Finally, we use AdamW [30] optimizer with a learning rate of and a batch size of 128 to fine-tune the pretrained LLaVA1.5 for one epoch which takes about 14 hours.
Training of Genixer. Here, we aim to train a data generator for grounding tasks such as region-based VQA that involve the object coordinates by referring to the corresponding objects. To ensure the generation, we utilize the entirety of the RefCOCO [20] and VG [21] datasets. Nonetheless, as indicated in Tab. 1, the dataset size for RD is relatively small. To counteract potential biases in the generation, Genixer that finetunes a SoTA model Shikra [9] on region-based tasks adopts two-phased training. The first phase focuses on the REC and REG data generation. In the second phase, Genixer adds PointQA, QCA, and RD data while deliberately reducing REC and REG data for data trade-off. To train the Genixer, we set the equal to 0.5 for each data type. We utilize the AdamW optimizer, applying learning rates of for the first phase and for the second phase. The batch sizes are set to 128 and 64, respectively.
3.4 Automatic Data Generation and Filtering
Here, we introduce how we use Genixer and Genixer to generate data and then present two effective data filtering strategies for Genixer and Genixer, respectively. This is important and necessary for ensuring the generated data quality and thus enhancing the performance.
Data Generation and Filtering of Genixer on The General Tasks. The diversity of image resources and effective data filtering are two key steps to obtaining high-quality instruction tuning data. Here, we first describe the image sources used for the data generation with Genixer. To prosper the diversity of current instruction tuning data, we utilize the mixed dataset comprising 558K images from the LAION [43], CC3M [44], and SBU [39], as described in [29], and further adopt the 830K images from the original SBU datasets. These different image resources can help to generate more diverse instruction tuning data. We directly feed these 1.4M images to the trained Genixer and then designate the specific instruction = “This is the Common VQA task”, and finally we produce the 1.4M raw VQA triplets pending for the next filtering.
To assess the quality of general task data, we design a Fuyu-driven data filtering framework to automatically and effectively filter the samples generated by Genixer. As illustrated in Fig. 6, we design the text prompt as follows:
This framework is applicable to filter the four general tasks defined in Sec. 3.1. To assess the data quality of Common VQA and Adv VQA tasks, we substitute the variables and with the generated questions and answers, respectively. In the case of MC VQA evaluation, we replace the option letter (e.g., “B”) with the corresponding option content (e.g., “Jumping”) and then convert the format to match that of Common VQA for processing by Fuyu-8B [5], as shown in Fig. 6. For the MD type, we decompose multi-turn dialogue into individual single-turn VQA instances, processing these in the same manner as Common VQA.
To avoid the potential bias to output “Yes” of Fuyu-8B, instead of directly getting the words Yes or No as the final results, we additionally calculate the probability of predicting the “Yes”. To this end, we restrict the output to either Yes or No and calculate the probability via the output logits as follows:
where is the predicted response and is the length the total response sequence. Then, we propose a threshold to control the filtering in the following manner:
where is the sign of keeping or removing the current sample, and denotes the probability of the result “Yes” of -th visual-question-answer triplet. By setting , we filter the 1.4M raw VQA triplets to 915K instances. We name this VQA-like synthetic dataset Genixer-915K.
Data Generation and Filtering of Genixer on The Grounding Tasks. Similarly, our Genixer adopts the same image resources in general tasks for generating grounding-based instruction tuning data. After feeding the 1.4M image corpus to Genixer by set the data type as REC, we get 1.4M raw data.
To assess the quality of these REC data, we propose a CLIP-driven data filtering framework. Specifically, we first use the regular expression to extract the text expression and region coordinates from the raw generated sentence and then conduct the following three steps to filter the generated data in a coarse-to-fine manner. 1) Removing the wrong question or answer formats (e.g., wrong coordinate format). 2) Removing the bonding box whose width or height is smaller than 50. 3) Employing OpenCLIP-L [19] model for calculating the similarity score between the text expression and their corresponding image region, discarding samples with CLIP scores below 0.6. For example, consider one REC sample with the Question “I need the coordinates of the person at the bottom left of the image. Can you assist?” and the Answer “[0.005,0.332, 0.249,0.984]”. Here, “person at the bottom left of the image” is the text expression, and the coordinate of the referring region is “[0.005,0.332,0.249,0.984]”.
By applying a threshold of 0.6, we filter out 350K instances from 1.4M images and name this REC-like synthetic dataset as Genixer-350K.
4 Experiments
In Sec. 3.4, we use two trained generators Genixer and Genixer to generate Genixer-915K and Genixer-350K for general and grounding tasks, respectively. Here, we evaluate the quality of these data from several aspects, including statistical analysis, human evaluation, evaluation via training MLLMs, and ablation study.
4.1 Statistical Analysis
To evaluate the generation quality of Genixer, we conduct a comparative analysis using the VQAv2 [15] training set as a comparison. Fig. 7 (a) illustrates the distribution of question lengths of VQAv2 and Genixer-915K. Notably, our dataset exhibits a significant long tail, indicating a higher proportion of longer sentences compared to VQAv2. Additionally, the distribution of nouns and verbs, depicted in Fig. 7 (b) and (c), showcases the diverse vocabulary present in the generated questions. To assess data quality, we utilize Flickr30K [52] as the image source and employ Genixer to generate three representative data types, as detailed in Sec. 3.4. Results in Tab. 2 demonstrate that our generated data achieves an accuracy exceeding 80%, with the highest accuracy observed in the MC VQA category. Furthermore, Eq. (3) guarantees that our generated data exhibits a high probability (0.8) of being classified as “Yes” across all three data types. This is corroborated by the probability distribution depicted in Fig. 9, affirming the high quality of data produced by Genixer in generating diverse instruction tuning data.
The statistics of the generated dataset Genixer-350K by Genixer are presented in Fig. 8, showcasing metrics related to expression length, region width, and height. Tab. 4 offers a comparative analysis of our dataset, highlighting the larger collection of images and expression lengths in Genixer-350K compared to other grounding-based datasets.
4.2 Human Evaluation
We conduct the human evaluation to manually analyze the generated question type and corresponding correctness. We employ Genixer to generate QA samples without specific the data type. Since the image content can affect the generation types, we randomly chose 100 images from the COCO validation set as the held-in set and 100 images from Flickr30K as the held-out set. As shown in Tab. 4, we divide the questions into seven types, such as “Action” and “Color”. One can observe that “Object Type” and “Relative position” are the two most popular question types of both held-in and held-out datasets. Among held-in samples, “Action” question also has a notable proportion with 92% correctness. Different from the held-in set, 38 out of 100 samples belong to “Object Type” and 23 out of 100 “Relative Position” in the held-out set. In all of these types, the correctness of held-in dataset is slightly higher than held-out dataset.
Question Type | Held-in (COCO val) | Held-out (Flickr30K) | ||
#Samples | Correct (%) | #Samples | Correct(%) | |
Action | 13 | 92 | 17 | 88 |
Color | 8 | 75 | 4 | 75 |
Counting | 6 | 83 | 3 | 66 |
Object Type | 23 | 87 | 38 | 76 |
Relative Position | 32 | 75 | 23 | 65 |
Yes/No | 2 | 50 | 4 | 100 |
Others | 16 | 81 | 11 | 82 |
4.3 Evaluation via Training MLLMs
Unless specifically stated otherwise, all experiments were conducted on an 8 Nvidia A100 (40G) GPU setup.
Evaluation on General Tasks. Considering the data size (665K) is small that is used for training LLaVA1.5 and the data balance during the finetuning stage, we proceed to add the Genixer-915K dataset into the pretraining phase, adhering to the same training protocols utilized by LLaVA1.5. Furthermore, we enriched the fine-tuning stage with an additional 8K VQA-like data points, selected based on Fuyu-8B’s probability range of 0.5 to 0.7. This probability range was chosen based on findings that higher probabilities correlate with simpler, more straightforward VQA instances. By selecting data within this specific range, our model is able to learn more challenging VQA samples. From Tab. 5, one can observe consistent enhancements across 10 out of 12 benchmarks compared with vanilla LLaVA1.5. Indeed, on several tasks, our generated data can make significant improvements, e.g., 3.8% on VizWiz, 2.9% on SciencQA, and 37.7 scores on the MME benchmark.
Method | LLM | VQA | GQA | VizWiz | SQA | VQA | POPE | MME | MMB | MMB | SEED | LLaVA | MM-Vet |
BLIP-2 [25] | Vicuna-13B | 41.0 | 41.3 | 19.6 | 61.0 | 42.5 | 85.3 | 1293.8 | – | – | 49.7 | 38.1 | 22.4 |
InstructBLIP [13] | Vicuna-7B | – | 49.2 | 34.5 | 60.5 | 50.1 | – | – | 36.0 | 23.7 | 58.8 | 60.9 | 26.2 |
InstructBLIP [13] | Vicuna-13B | – | 49.5 | 33.4 | 63.1 | 50.7 | 78.9 | 1212.8 | – | – | – | 58.2 | 25.6 |
Shikra [9] | Vicuna-13B | 77.4 | – | – | – | – | – | – | 58.8 | – | – | – | – |
IDEFICS-9B [22] | LLaMA-7B | 50.9 | 38.4 | 35.5 | 44.2 | 25.9 | – | – | 48.2 | 25.2 | 44.5 | – | – |
IDEFICS-80B [22] | LLaMA-65B | 60.0 | 45.2 | 36.0 | 68.9 | 30.9 | – | – | 54.5 | 38.1 | 53.2 | – | – |
Qwen-VL [3] | Qwen-7B | 78.8 | 59.3 | 35.2 | 67.1 | 63.8 | – | – | 38.2 | 7.4 | 62.3 | – | – |
Qwen-VL-Chat [3] | Qwen-7B | 78.2 | 57.5 | 38.9 | 68.2 | 61.5 | – | 1487.5 | 60.6 | 56.7 | 65.4 | – | – |
LLaVA-1.5 | Vicuna-7B | 78.5 | 62.0 | 50.0 | 66.8 | 58.2 | 85.9 | 1465.0 | 64.3 | 58.3 | 66.2 | 65.4 | 31.1 |
LLaVA-1.5 | Vicuna-7B | 79.1 | 63.1 | 53.8 | 69.7 | 59.0 | 87.3 | 1502.7 | 65.3 | 59.4 | 66.6 | 64.0 | 30.1 |
Method | RefCOCO | RefCOCO+ | RefCOCOg | Avg | |||||
val | test-A | test-B | val | test-A | test-B | val | test | ||
OFA-L* [48] | 79.96 | 83.67 | 76.39 | 68.29 | 76.00 | 61.75 | 67.57 | 67.58 | 72.65 |
OFASys [4] | 80.10 | – | – | – | – | – | – | – | – |
VisionLLM-H [50] | 86.70 | – | – | – | – | – | – | – | – |
Shikra [9] | 87.01 | 90.61 | 80.24 | 81.60 | 87.36 | 72.12 | 82.27 | 82.19 | 82.92 |
Shikra | 87.48 | 91.05 | 81.77 | 81.89 | 87.43 | 73.14 | 81.99 | 83.15 | 83.49 |
Evaluation on Grounding Tasks. Here, we adopt Shikra [9] as the model to evaluate the quality of Genixer-350K REC-like data produced by Genixer. Adhering to Shikra’s published code, we incorporate our synthetic data into the training phases, maintaining consistent training iterations to ensure a fair comparison. Tab. 6 shows the improvement on 7 out of 8 test datasets with a non-trivial average boost of 0.6%. These findings imply that our pipeline can be an alternative approach to generate grounding-based instruction tuning data, which is typically challenging for manually labeling and not satisfied for prompting GPT-4V, as shown in Fig. 2.
Performance on Genixer. As illustrated in Tab. 2 and Fig. 9, Genixer showcases a superior capability of generating high-quality data. It is natural for us to investigate the performance of Genixer. Accordingly, we evaluate Genixer and report its results in Tab. 7, which refers to the setting ZS. One can observe the minor declines in performance on the GQA, ScienceQA, and SEED, alongside a modest enhancement on VQAv2, VizWiz, and POPE. Such results are due to the exclusive training on generating instruction tuning data. Thus, for a fair comparison to investigate the capability of Genixer, we proceeded to retrain the Genixer with the mixture of the 665K finetuning dataset used in LLaVA1.5 and the datasets for training Genixer for one epoch following the same training protocols. The outcomes of this process are presented as MixT in Tab. 7, where we witnessed significant improvements across all six benchmarks.
Method |
Setting |
VQAv2 | GQA | VizWiz | SQA | POPE | SEED |
LLaVA1.5 |
78.5 | 62.0 | 50.0 | 66.8 | 85.9 | 66.2 |
Genixer |
ZS |
79.7 | 61.3 | 50.8 | 65.6 | 86.0 | 65.6 |
Genixer |
MixT |
80.2 | 63.1 | 54.1 | 67.1 | 87.5 | 67.1 |
Dataset | VQAv2 | GQA | VizWiz | SQA | POPE | SEED |
Baseline | 78.5 | 62.0 | 50.0 | 66.8 | 85.9 | 66.2 |
Genxier-300K | 79.0 | 62.9 | 52.7 | 68.5 | 87.1 | 65.8 |
Genxier-610K | 79.0 | 63.1 | 53.7 | 69.2 | 87.2 | 66.2 |
Genixer-915K | 79.1 | 63.1 | 53.8 | 69.7 | 87.3 | 66.6 |
Setting | Size | VQAv2 | GQA | VizWiz | SQA | POPE | SEED |
Baseline | - | 78.5 | 62.0 | 50.0 | 66.8 | 85.9 | 66.2 |
1.4M | 79.0 | 62.9 | 53.5 | 69.6 | 87.1 | 66.2 | |
1.1M | 79.1 | 63.1 | 53.2 | 69.1 | 86.9 | 66.4 | |
0.9M | 79.1 | 63.1 | 53.8 | 69.7 | 87.3 | 66.6 |
4.4 Ablation Study
Effect of data scale. Tab. 9 investigates the effects of the scales of our synthetic data in the pretraining stage. One can observe that a larger scale often leads to a steady performance improvement on all of the six benchmarks, showing the quality of our synthetic data.
Probability Range. Tab. 9 investigates the impact of the probability threshold during data filtering in Section 3.4 on the data quality. By varying , we observe that higher values of often better improve performance across all six benchmarks, even with a reduced number of selected training samples. This suggests that the quality of data is more crucial than the quantity of samples.
5 Conclusion, Limitations, and Societal Impacts
In this paper, we introduce a novel automatic data generation pipeline called Genixer, designed to efficiently and affordably produce high-quality instruction tuning data by leveraging current MLLMs. We instantiate Genixer into two data-generative MLLMs, Genixer and Genixer, tailored to generate general and grounding instruction tuning data, respectively. To ensure the quality of the generated data, we propose two data filtering frameworks: Fuyu-driven and CLIP-driven. Finally, we contribute two instruction tuning datasets, Genixer-915K and Genixer-350K, targeting Common VQA and REC. Experimental results demonstrate that both generated datasets significantly enhance LLaVA1.5 and Shikra across various multimodal benchmarks, respectively.
Limitations. 1) LLM Scale: Due to computational constraints, we do not test larger LLM model, such as 13B or 34B. But we believe that our data generator could be beneficial, since larger models are more data hungry. 2) Data Scale: While scaling up the candidate image corpus to larger datasets like LAION-2B could enhance model capability, training costs and time constraints restrict us to do such expansions. But Tab. 9 shows scaling can improve performance. 3) Evaluation: Despite proposing effective data filtering frameworks, evaluating complex and open-ended data types like Referential Dialogue remains challenging, leaving room for future exploration.
Societal Impacts. Our work addresses the challenge of generating high-quality instruction tuning data by presenting a comprehensive pipeline. It paves the way for future investigations into generating diverse multimodal data, contributing to advancements in various fields.